Circular No. 6285

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J1744–28 M. H. Finger, Compton Observatory Science Support Center; R. B.

Wilson and B. A. Harmon, Marshall Space Flight Center, NASA; K. Hagedon, University Space Research Association; and T. A. Prince, California Institute of Technology, report for the Compton Observatory BATSE team: “Pulsations with a period of 467 ms are being detected in the energy band 25–45 keV from a source in the galactic-center region. The pulsed flux is ∼ 40 percent of the total flux observed from the hard x-ray transient GRO J1744–28 (IAUC 6284). Absence of any other persistent source consistent with the direction and flux of the observed pulsed emission implies that the pulsations are from GRO J1744–28. The 31-ms-resolution data in which the pulsations were observed excluded time intervals containing outbursts from the possibly separate burst source in the same region (IAUC 6272, 6275, 6276). On Jan. 6.0 UT, the measured barycentric pulse frequency was 2.1409720(5) Hz.”

OB AND X-RAY BINARY CANDIDATES C. Motch and M. Pakull, CNRS, Observatoire Astronomique de Stras-

bourg; F. Haberl and K. Dennerl, Max-Planck-Institut für Extraterrestrische Physik, report the discovery of several new galactic OB/x-ray binary candidates: “These massive systems were selected from a cross-correlation of the ROSAT all-sky survey (|b| < 20^{o}) with OB star catalogues, and display significant excess x-ray emission (range 0.1–2.4 keV) over the expected stellar level. Follow-up optical and ROSAT observations yield four very likely new massive x-ray binaries: the O star LS 5039 (RX J1826.2–1450) and the Be stars BSD 24-491 (RX J0440.9+4431), LS 992 (RX J0812.4– 3114) and LS 1698 (RX J1037.5–5647). LS 992 exhibited an x-ray outburst during survey observations (1990 Oct. 28.5–31.1 UT; x-ray luminosity 1.3 × 10^{35} erg/s, assuming d = 9.2 kpc), and LS 1698 is probably identical to the hard x-ray transient 4U 1036–56. The new candidates have x-ray luminosity > 6 × 10^{33} erg/s, indicating the presence of an accreting neutron star or black hole. Two further Be/x-ray binary candidates, HD 161103 and SAO 49725 require confirmation of their x-ray excess (x-ray luminosity 1–5 × 10^{32} erg/s) and could be Be + accreting white dwarf systems.”

1996 January 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green